OpenStudy (anonymous):

determine whether (-9, -5) is a solution of 6x+7y=20

OpenStudy (anonymous):

Just plug the values in. x=-9, y=-5 6*(-9)+7*(-5)=20 -89 = 20 => false Therefore (-9, -5) is not a solution.